OpenStudy (anonymous):

A racing dog is initially running at 13.0 , but is slowing down.How fast is the dog moving when its kinetic energy has been reduced by half? By what fraction has its kinetic energy been reduced when its speed has been reduced by half?

OpenStudy (anonymous):

Initial Kinetic energy = Ki initial velocity = Vi = 13.0 m/s Slower velocity = Vs Slower Kinetic Energy = Ks general kinetic energy \[K=\frac{ 1 }{ 2 }mv^{2}\] general velocity \[v^{2}=\frac{ 2K }{ m }\] A) \[K_{i}=\frac{ 1 }{ 2 }mv_{i}^{2}\] B) \[v_{s}^{2}=\frac{ 2\frac{ K_{i} }{ 2 } }{ m }\] C) \[K_{f}=\frac{ 1 }{ 2 }m(\frac{ v_{i} }{ 2 })^{2}\]

OpenStudy (anonymous):

derp, B = A, C = B.